# Batayan: A Filipino NLP benchmark for evaluating Large Language Models

### v1

@@ 4 Challenges with Developing a Filipino Benchmark
This section discusses issues in creating Batayan, split into datasets that required adaptation and datasets where the content were kept as-is. The purpose of this section is to shed light on issues that future researchers may encounter in creating new Filipino datasets or adapting previous ones.
4.1 Issues with Adapted Data
Most of the tasks presented in this study were sourced from existing English language datasets that were natively translated and adapted into Filipino by the authors, whom are all native speakers of Filipino.
@@ 4.1 Issues with Adapted Data
Limitations in adapting vocabulary. Adapting the English tasks to Filipino revealed issues in maintaining the intelligibility of the text. Technical terms that exist in English were difficult to translate into Filipino, with several of them having no direct translations. For example, an instance from Belebele included the term “rule of thirds”, which describes a specific photography technique. Originally, the translation of this phrase in the Tagalog subset of Belebele was “tuntunin ng mga sangkatlo” (lit. “rule of thirds”), which is does not make sense in Filipino and does not convey the intended meaning. For jargon such as this, we chose to keep the original English phrasing.
English idiomatic expressions were also prevalent and required a more flexible approach in adaptation to ensure the integrity of meaning. For example, one premise from Balanced COPA was, “The criminal turned himself in”, using the idiomatic expression “turned himself in” to mean “surrendered”. We chose to translate this into Filipino as, “Isinuko ng kriminal ang kanyang sarili” (in English: “The criminal surrendered himself”).
Moreover, the presence of English words that have homonyms added to the complexity of translating English sentences to Filipino. For example, some sentences in the PAWS dataset contained the adjective “right”, which can refer to either the direction (in Filipino: “kanan”) or an assertion of the correctness of the object it is modifying (in Filipino: “tama”). In such cases, we inferred the most probable meaning of the words from the context of the entire passage and then identified the corresponding Filipino word to be incorporated in our Filipino translation.
(Dis)fluency in expert and machine translations. Our rigorous review and annotation of existing datasets with machine and expert-guided translations in Filipino revealed weaknesses in their adaptation. The authors found that the automatic translations were unnatural, disfluent, and showed characteristics of translationese Gellerstam (1986); Riley et al. (2020) despite being grammatically correct. Notably, we observed that these initial translations demonstrated an unusual preference with using ay-inversion. We also observed that these translations used Filipino terms that were synonymous to their English counterparts but pragmatically-awkward. Hence, we applied rephrasing and re-translation to ensure that the samples used sound more native and natural.
@@
Inconsistencies in ay-inversions in translation construction. Styles in translation vary from person-to-person, depending on their valuations and reading of the original text Castagnoli (2020). Likewise, ay-inversion in the context of translation boils down to a stylistic choice of the translator. Hence, while the KA is deemed more natural over the DKA, this valuation of the usual/natural and unusual/unnatural is debated.
On one hand, university stylebooks on writing such as that of Yapan (2017) considers KA to mean the usual pattern of speech for Filipino speakers. Also, Magracia (2001) notes that while the DKA is correct in “meaning, syntactic order, and grammaticality,” it is not the “natural” way of speaking for Filipinos. Furthermore, she argues that the DKA reflects a speaker’s language of thought—in this case, the influence of English. The same justification, in terms of source language for translated corpora as demonstrated in Visweswariah et al. (2011), can then be said on the machine-translated text, though this remains an open question for Filipino.
On the other hand, other research argues that DKA is not an abnormality, but is instead a necessary aspect for “discourse continuity” Bolata (2022). Hence, the more complex the sentence, the higher the chance of utilizing the ay-inversion Fox (1985). The same can be observed in the datasets. For example, the short sentences in the CR task allowed for the sentences to be changed to the KA, while longer, more context-rich sentences in the NLI task required the use of the DKA.
These frameworks demonstrate the subjective nature of the selection of sentence structure. Hence, such inconsistencies may still be observed in the tasks in Batayan despite being rigorously translated by native speakers.
Models PI QA SA TD CR NLI
aisingapore/gemma2-9b-cpt-sea-lionv3-instruct 82.02 82.80 75.99 73.12 92.75 68.31
aisingapore/llama3.1-8b-cpt-sea-lionv3-instruct 74.98 80.69 51.18 61.24 79.70 66.26
CohereForAI/aya-expanse-8b 33.94 2.97 57.44 34.39 1.96 25.27
google/gemma-2-9b-it 82.50 83.96 70.72 64.26 91.25 60.61
meta-llama/Llama-3.1-8B-Instruct 60.94 73.97 52.54 49.76 73.21 57.64
Qwen/Qwen2.5-7B-Instruct 84.45 69.11 67.42 61.05 45.76 58.75
sail/Sailor2-8B-Chat 48.93 59.48 48.40 47.48 19.82 24.05
SeaLLMs/SeaLLMs-v3-7B-Chat 36.98 49.68 70.79 63.26 71.91 20.07
Table 2: Performance of LLMs on Batayan natural language understanding (NLU) and natural language reasoning (NLR) tasks. Macro F1 scores are reported.
@@ 4.2 Issues with Natively-Sourced Data
Non-standard orthography. The dataset exhibited significant orthographic variation (spelling, capitalization, word boundaries, etc.) due to its natural language origins. We preserved this variation as it reflects authentic Filipino language Ilao et al. (2012); Javier (2018); Caroro et al. (2020) use relevant to toxicity detection and sentiment analysis tasks, aligning with our methodological principles. We maintained quality control by verifying that orthographical variations remained comprehensible to native readers.
Class imbalance. The dataset showed a disproportionate distribution of negative sentiment regarding specific political entities. We addressed this by rebalancing the dataset and prioritizing high-agreement samples to mitigate individual annotator bias.
@@ 5 Evaluation
Batayan serves as the Filipino component of the SEA-HELM leaderboard. The evaluation framework of SEA-HELM follows prior systems such as HELM Liang et al. (2023), and is composed of tasks, prompts, and metrics.
@@ 5.1 Evaluation Design
Tasks. The selection of tasks in Batayan aforementioned in Section 3.1 aims to comprehensively evaluate language capabilities of LLMs in Filipino. Each task is a collection of instances composed of an input string, a list of references, accompanying metadata, and the groundtruth label. We divide instances into two sets: the evaluation set, and a small 5-instance set that can serve as exemplars for few-shot prompting. Overall, Batayan provides 8 distinct tasks with 3,800 test instances.
Prompts. For Batayan, we developed prompt templates for each task written in Filipino. We ensured that the instructions for each task are consistent with the prompt template design already used in SEA-HELM. A comprehensive list of these prompt templates can be found in Appendix C.
During model evaluation, input prompts are constructed using evaluation instances and the corresponding prompt template. The default evaluation setting for Batayan is zero-shot prompting, where in-context input-label pairs are not included in the model prompt. Given the input prompts and decoding parameters (see Appendix LABEL:sec:experimental_setup), a model then generates output completions.
@@ 5.2 Results
Models AS MT (eng→tgl) MT (tgl→eng)
BERTScore ChrF++ ROUGE-L F1 ChrF++ MetricX-24 ChrF++ MetricX-24
aisingapore/gemma2-9b-cpt-sea-lionv3-instruct 73.08 34.43 21.06 57.80 87.23 65.93 90.55
aisingapore/llama3.1-8b-cpt-sea-lionv3-instruct 73.47 34.15 23.82 57.08 82.46 62.73 87.13
CohereForAI/aya-expanse-8b 59.38 6.38 1.39 39.27 37.47 59.23 78.88
google/gemma-2-9b-it 72.68 34.09 21.73 56.64 84.00 67.54 76.49
meta-llama/Llama-3.1-8B-Instruct 74.33 34.51 27.39 44.91 51.92 54.28 72.04
Qwen/Qwen2.5-7B-Instruct 70.50 28.74 15.77 38.78 34.23 54.94 75.97
sail/Sailor2-8B-Chat 71.84 32.06 19.82 50.44 79.85 53.34 20.12
SeaLLMs/SeaLLMs-v3-7B-Chat 70.55 28.23 16.90 46.03 61.29 39.03 59.97
Table 3: Performance of LLMs on Batayan natural language generation (NLG) tasks.
[…]
Analysis of model performance across tasks reveals several insights about current multilingual LLMs’ capabilities in Filipino. Models with explicit Filipino language support generally outperform those without dedicated instruction tuning, though the margin varies significantly by task type. In NLU and NLR tasks (see Table 2), gemma2-9b-cpt-sea-lionv3-instruct demonstrates superior performance, achieving the highest scores in SA (75.99%), TD (73.12%), and NLI (68.31%). This suggests that targeted regional pretraining and instruction tuning can capture the nuances of Filipino sentiment and toxicity, which often involve code-switching and cultural context. However, for PI, Qwen2.5-7B-Instruct achieves the best performance (84.45%), indicating that semantic similarity tasks may benefit more from general multilingual pretraining than from region-specific tuning.
The performance gap between models with Filipino language support and general multilingual models is most pronounced in CR, where gemma2-9b-cpt-sea-lionv3-instruct (92.75%) significantly outperforms models without dedicated Filipino support. This suggests that understanding causality in Filipino requires strong grasp of the language’s complex morphological system, particularly in how verb affixes encode causative relationships.
In MT (see Table 3), Filipino-supported LLMs show stronger performance overall. For English→Filipino translation, gemma2-9b-cpt-sea-lionv3-instruct leads in both ChrF++ (57.80) and MetricX-24 (87.23). However, for Filipino→English translation, while gemma-2-9b-it achieves a slightly higher ChrF++ (67.54) than gemma2-9b-cpt-sea-lionv3-instruct (65.93), it is substantially overtaken by the latter on MetricX-24 (90.55 vs. 76.49). This performance gap suggests that region-specific tuning not only aids comprehension of Filipino source text but can also the semantic fidelity of English outputs—an aspect captured more effectively by MetricX-24.
In AS, Llama-3.1-8B-Instruct achieves the highest BERTScore (74.33), despite lacking explicit Filipino support. This unexpected result may indicate that general abstraction and summarization capabilities transfer well across languages, even without language-specific tuning. However, when examining ROUGE-L scores, Filipino-supported models show stronger performance, suggesting they better preserve Filipino-specific discourse structures and coherence patterns.
In our experiments, we found that the aya-expanse-8b model mostly failed to handle Filipino instructions, resulting in low scores particularly in CR (1.96%) and QA (2.97%).
Examining the relationship between model size and performance reveals an interesting pattern: while larger models generally perform better, the correlation is weaker for Filipino-specific tasks compared to similar evaluations in English. This suggests that current scaling approaches may not adequately capture the unique characteristics of Filipino, such as its agglutinative morphology and frequent code-switching.
@@ 6 Conclusion
In this paper, we introduced Batayan for holistically evaluating LLMs on a gamut of Filipino language tasks covering natural language understanding, reasoning, and generation. Our findings show that LLMs with explicit Filipino language support and finetuned on Filipino instructions demonstrate better performance on Batayan compared to other models.
@@
This project was approved by the principal investigator’s university internal review board.
@@
For the review and annotation of the sentiment analysis task, we involved a quality assurance team consisting of native Filipino speakers. The team, comprised of students at local universities, were recruited through public advertisements that stated the estimated work load and remuneration, which were consistent with university research guidelines and regulatory requirements.
@@
We plan to release Batayan (datasets and codebase) under the Creative Commons Attribution Share-Alike 4.0 (CC-BY-SA 4.0) license.
@@ Acknowledgments
This research project is supported by the National Research Foundation, Singapore under its AI Singapore’s National Large Language Models Funding Initiative.
@@
Appendix A Inter-rater Agreement

### v2

@@ 4 Challenges with Developing a Filipino Benchmark
This section discusses issues in creating Batayan, shedding light on issues that future researchers may encounter in creating new Filipino datasets or adapting previous ones.
4.1 Issues with English-adapted and English-sourced Data
@@
Limitations in adapting vocabulary. Adapting the English tasks to Filipino revealed issues in maintaining the intelligibility of the text. Technical terms that exist in English were difficult to translate into Filipino, with several of them having no direct translations. For example, an instance from Belebele included the term “rule of thirds”, which describes a specific photography technique. Originally, the translation of this phrase in the Tagalog subset of Belebele was “tuntunin ng mga sangkatlo” (lit. “rule of thirds”), which does not make sense in Filipino and does not convey the intended meaning. For jargon such as this, we chose to keep the original English phrasing.
English idiomatic expressions were also prevalent and required a more flexible approach in adaptation to ensure the integrity of meaning. For example, one premise from Balanced COPA was, “The criminal turned himself in”, using the idiomatic expression “turned himself in” to mean “surrendered”. We chose to translate this into Filipino as, “Isinuko ng kriminal ang kanyang sarili” (lit.: “The criminal surrendered himself”).
Moreover, the presence of English words that have homonyms added to the complexity of translating English sentences to Filipino. For example, some sentences from PAWS contained the adjective “right”, which can refer to either the direction (in Filipino: “kanan”) or an assertion of the correctness of the object it is modifying (in Filipino: “tama”). In such cases, we inferred the most probable meaning of the words from the context of the entire passage and then identified the corresponding Filipino word to be incorporated in our Filipino translation.
(Dis)fluency in expert and machine translations. Our rigorous review of existing English datasets with machine and expert-guided translations in Filipino revealed weaknesses in their adaptation. The authors found that the automatic translations were unnatural, disfluent, and showed characteristics of translationese despite being grammatically correct. Notably, we observed that these initial translations demonstrated an unusual preference with using ay-inversion, and used Filipino terms that were synonymous to their English counterparts but pragmatically-awkward. Hence, we applied re-translation to ensure that the samples sounded more native and natural.
@@
Inconsistencies in ay-inversions in translation construction. Styles in translation vary from person-to-person, depending on their valuations and reading of the original text Castagnoli (2020). Likewise, ay-inversion in the context of translation boils down to a stylistic choice of the translator.
While KA is argued to be the natural pattern of speech for Filipino speakers Yapan (2017), this valuation of natural and unnatural is debated. Magracia (2001) notes that while DKA is correct in “meaning, syntactic order, and grammaticality,” it is not the “natural” way of speaking for Filipinos, as it reflects the speaker’s language of thought—in this case, the influence of English. The same justification, in terms of source language for translated corpora, can be said for machine-translated text Visweswariah et al. (2011), though this remains an open question for Filipino.
On the other hand, other research argues that DKA is a necessary aspect for “discourse continuity” Bolata (2022). Hence, the more complex the sentence, the higher the chance of utilizing the ay-inversion Fox (1985). We observed this in some datasets such as NLI, which had longer, more context-rich sentences that required the use of DKA in contrast with the shorter, KA-ordered sentences found in the CR task.
These frameworks demonstrate the subjective nature of the selection of Filipino sentence structure. Hence, such inconsistencies may still be observed in Batayan despite being rigorously translated by native speakers.
@@ 4.2 Issues with Natively-Sourced Data
Non-standard orthography. The dataset exhibited significant orthographic variation (spelling, capitalization, word boundaries, etc.) due to its natural language origins. We preserved this variation as it reflects authentic Filipino language use Ilao et al. (2012); Javier (2018); Caroro et al. (2020) relevant to toxicity detection and sentiment analysis tasks, aligning with our methodological principles. We maintained quality control by verifying that orthographical variations remained comprehensible to native readers.
Class imbalance. The dataset showed a disproportionate distribution of negative sentiment for certain political entities. We rebalanced the class distribution and prioritized high-agreement samples to mitigate individual annotator bias.
@@ 5 Evaluation
Batayan serves as the Filipino component of the SEA-HELM leaderboard. The evaluation framework of SEA-HELM follows prior systems such as HELM Liang et al. (2023).
@@ 5.1 Evaluation Design
Tasks. The selection of tasks in Batayan mentioned in Section 3 aims to comprehensively evaluate the Filipino language capabilities of LLMs. Each task is a collection of test and example instances composed of an input string, references, metadata, and a ground truth label.
Prompts. For Batayan, we developed prompt templates for each task written in Filipino. We ensured that the instructions for each task are consistent with the prompt template design already used in SEA-HELM. A comprehensive list of these prompt templates can be found in Appendix D.
During model evaluation, input prompts are constructed using evaluation instances and the corresponding prompt template. The default evaluation setting of Batayan for instruction-tuned models is zero-shot prompting, where model prompts do not include in-context input-label examples. Base pre-trained models without instruction-tuning are evaluated under a five-shot setting. Given the input prompts and decoding parameters (see Appendix LABEL:sec:experimental_setup), a model then generates output completions.
@@ 5.2 Results
Figure 3: Model performance of selected models on NLU, NLR, and NLG tasks. NLU and NLR tasks are measured using macro F1 score, NLG tasks are measured using ChrF++, MetricX-24, BERTScore, and ROUGE-L F1. The unweighted average across tasks and scores is taken. Targeted Filipino instruction tuning allows Gemma-SEA-LION-v3-9B-IT to achieve competitive performance with significantly larger open-source and commercial models.
Our evaluation results include over 50 small ($\leq$11B), medium ($\leq$32B), and large ($\geq$32B) LLMs, including several commercial AI systems on the Batayan benchmark. The results underscore two principal findings: the substantial benefit of explicit Filipino language support and the scaling effects observed across model sizes. Figure 3 highlights several model performances, and the complete experiment results can be found in Appendix LABEL:sec:complete_experimental_results.
Filipino language support enhances task-specific performance. Tables LABEL:tab:small_nlu_nlrLABEL:tab:commercial_nlu_nlr in Appendix LABEL:sec:complete_experimental_results reveal that models with Filipino instruction tuning (such as the SEA-LION family of models) consistently outperform counterparts on culturally nuanced tasks. In particular, the Filipino-adapted model Gemma-SEA-LION-v3-9B-IT achieves an average macro F1 of 79.23 among small models, with a striking CR score of 92.75. In contrast, non-Filipino-specific models, such as many variants in the Meta Llama 3 series, exhibit very low or zero performance in CR, which emphasizes the challenge of capturing Filipino’s complex causality and morphology without explicit tuning.
On the NLG competency, tables LABEL:tab:small_nlgLABEL:tab:commercial_nlg in Appendix LABEL:sec:complete_experimental_results show that while ROUGE-L F1 scores show only modest differences between Filipino-tuned and general models, complementary evaluation using MetricX-24 provides a more sensitive measure of semantic fidelity. For instance, in engtgl translation, Filipino-supported systems like Gemma-SEA-LION-v3-9B-IT and Sailor2-8B-Chat achieve MetricX-24 scores exceeding 87, indicating superior preservation of Filipino-specific discourse and semantics. This observation shows that the apparent parity in ROUGE-L is only in part: when examined alongside semantic metrics, the benefit of dedicated language tuning is clear.
Scaling effects have clear trends beyond parameter similarity. Although many small models lie within a narrow 7B–9B parameter range, our broader evaluation—including medium and large models—illustrates a clear scaling effect. For example, among large models, Gemma-SEA-LION-v3-9B-IT attains an average macro F1 of 78.41 for NLU, macro F1 of 80.69 for NLR, and aggregated score of 60.35 in NLG, significantly outperforming similarly-tuned small models. This trend is evident even when models are compared within the same family: larger variants consistently demonstrate enhanced capability in handling both the nuanced understanding and generation challenges posed by Filipino text. The results thus suggest that increased model capacity—when combined with region-specific instruction tuning—enables a better representation of Filipino linguistic phenomena, including code-switching and agglutinative morphology.
Insights from commercial systems. In addition to open-source models, our evaluation also includes several commercial LLMs (Tables LABEL:tab:commercial_nlu_nlr and LABEL:tab:commercial_nlg). Commercial systems such as Gemini and GPT-4o variants exhibit competitive performance across both understanding and generation tasks. For instance, on NLU and NLR tasks, commercial models achieve average macro F1 scores ranging from 75.14 to 86.23. On the NLG competency, commercial systems also deliver robust results: Gemini and GPT-4o models maintain strong MetricX-24 scores above 88, indicative of high-quality semantic translation in both engtgl and tgleng directions.
While these commercial models benefit from large-scale pre-training and domain-specific optimization, our findings show that open-source, Filipino-tuned models achieve comparable—and in certain tasks, superior—performance. In particular, when cultural and linguistic nuances are critical, explicit regional instruction tuning offers a tangible advantage. The comparable performance of commercial systems reinforces that our dual strategy of targeted language support combined with model scaling is effective in academic settings as well as real-world applications.
Overall, our results provide strong evidence for a dual strategy in multilingual modeling: integrating region-specific instruction tuning with scalable architectures. Filipino-tuned models excel in tasks that require nuanced cultural and morphological sensitivity, and their performance improves markedly with increased capacity. Notably, open-source models fine-tuned with Filipino instructions are competitive with and even surpass state-of-the-art commercial systems on both understanding and reasoning tasks. These findings highlight that explicit Filipino language support not only enhances semantic fidelity and discourse preservation but also positions such models ideally for real-world, production-level applications in the Philippines.
@@ 6 Conclusion
In this paper, we introduced Batayan for holistically evaluating LLMs on a gamut of Filipino language tasks covering natural language understanding, reasoning, and generation. Our findings show that LLMs with explicit Filipino language support and fine-tuned on Filipino instructions demonstrate better performance on Batayan compared to other models.
@@
While we recognize that datasets with larger amounts of samples are preferred to ensure generalizability, we maintain our position to prioritize consistent, high-quality, and representative samples. The thoroughness and rigor of our review, annotation, and translation processes were essential and maintaining the quality and reliability of our benchmark. We believe that our benchmark serves as a valuable starting point for future research and can certainly be expanded in subsequent iterations.
Further, given the breadth of models and tasks, a comprehensive evaluation incorporating error cases was not possible. We also recognize the importance of evaluation LLMs that specialize on specific tasks such as translation. We hope that a public release of the Batayan codebase and benchmark on a leaderboard can help researchers better characterize the weaknesses of models, conduct their own analysis on the Filipino language capabilities of LLMs, and experiment with specialized models on Batayan.
@@
The methodology used in this work was approved by an Institutional Review Board (NUS-IRB-2024-617). We have released the Batayan dataset as part of the SEA-HELM evaluation suite and leaderboard under the Creative Commons Attribution Share-Alike 4.0 (CC-BY-SA 4.0) license, respecting the licenses of the source datasets used in this study.
@@
For the annotation of the sentiment analysis task, we involved a quality assurance team consisting of native Filipino speakers. The team, comprised of students from local universities in Singapore, were recruited through public advertisements. These stated the estimated work load and remuneration, which were consistent with university research guidelines and regulatory requirements.
@@ Acknowledgments
This research/project is supported by the National Research Foundation, Singapore under its National Large Language Models Funding Initiative. Any opinions, findings and conclusions or recommendations expressed in this material are those of the authors and do not reflect the views of National Research Foundation, Singapore.
@@
Appendix A Overview of Tasks and Datasets
Competency Task Dataset License
NLU PI PAWS Zhang et al. (2019) CC BY 4.0
QA Belebele Bandarkar et al. (2024) CC BY-NC 4.0
SA PH Elections Cabasag et al. (2019) Unknown
TD PH Elections Cabasag et al. (2019) Unknown
NLR CR Balanced COPA Kavumba et al. (2019) CC BY 4.0
NLI XNLI Conneau et al. (2018) CC BY-NC 4.0
NLG AS XL-Sum Hasan et al. (2021) CC BY-NC-SA 4.0
MT FLORES 200 NLLB Team et al. (2022) CC BY-SA 4.0
Table 3: License details for datasets used in Batayan.
Competency Task # test samples Avg. No. Words/Sample
NLU PI 400 22.50 (Sentence 1), 22.47 (Sentence 2)
QA 100 14.75 (Question), 4.26 (Choices)
SA 600 24.32
TD 400 20.73
NLR CR 400 7.24 (Premise), 6.05 (Choices)
NLI 600 23.30 (Sentence 1), 12.20 (Sentence 2)
NLG AS 100 120.8
MT (English Text) 600 21.42
MT (Tagalog Text) 600 25.05
Table 4 presents quantitative statistics on dataset size and average word counts. The variation in text length reflects task-specific requirements: AS involves longer passages, while classification tasks (SA, TD, and PI) typically contain more concise text samples. The QA dataset, in particular, has relatively short questions and answer choices, mirroring real-world multiple-choice assessments. Further granularity is provided in Table 5, which breaks down sentence, word, and character-level statistics. This statistics shows structural differences between tasks; for instance, MT has sentence-level parallelism, while CR and NLI involve distinct premise-hypothesis or question-choice relationships.
NLU PI Sentence 1 1.0 22.50 123.01
PI Sentence 2 1.0 22.47 122.99
QA Question 1.0 14.75 80.62
QA Choice 1 1.0 4.66 29.10
QA Choice 2 1.0 4.61 28.60
QA Choice 3 1.0 4.76 29.74
QA Choice 4 1.0 3.0 28.58
SA Text 2.21 24.32 120.44
TD Text 2.45 20.73 99.37
NLR CR Premise 1.0 7.24 37.55
CR Choice 1 1.0 6.05 30.91
CR Choice 2 1.0 6.05 30.62
NLI Sentence 1 1.0 23.30 127.88
NLI Sentence 2 1.0 12.20 67.04
NLG AS Text 5.43 120.8 666.18
MT English Text 1.12 21.42 113.07
MT Filipino Text 1.12 25.05 140.58
Appendix B Dataset Examples
Competency Task Example
Label: Paraprase (Paraphrase)
Question: Ano ang pinaniniwalaang dahilan ng pagbagsak?
Label: 1
Label: Negatibo (Negative)
Label: Malinis (Clean)
NLR CR Text: Hindi tinanggap ang tsekeng ginawa ko.
Question: Sanhi (Cause)
Label: 0
Sentence 2: Hindi kami pupunta sa birthday party ni Tita Basia ngayon.
Label: Contradiction
Table 6: Example instances for each task in Batayan.
Appendix C Inter-rater Agreement